the percentage change in real GDP the percentage change in nominal investment the percentage change in real investment not all economic activities are reported to the government. O people are not always paid as much as they are worth. 120 125 116 125 19%
Answer #1

Answer : Option B is correct. When economics talk about the economic growth of the country then it means that they are talking about the change in the real GDP it means that remain constant and the quantity change has been taken into account.

Answer 11: Option C is correct. As significant share of GDP is not included in the in the governments that's because all the economic activities are not reported to the government so there is gap between the actual GDP and statistical information.

Answer 12 : Option B is correct.

CPI of year 1 = ( Price of goods in year 1/ Price of goods in base year)*100= ( 750/625)*100= 120

Answer 13 : Option C is correct.

GDP Deflator = ( Nominal GDP/Real GDP)*100=( 850/735)*100= 116

Answer 14 : In this base year CPI = 100

than CPI in year 1= 119

Inflation rate = ((119-100)/100)*100= 19%